# On Disentangled Training for Nonlinear Transform in Learned Image Compression

### v1

@@ 1 Introduction
In fact, efficient training of LIC models has not been thoroughly explored, despite the properties of nonlinear transforms have been studied. Previous studies (Duan et al., 2022; He et al., 2022; Li et al., 2023) find that the analysis transform in LIC exhibits energy compaction property (i.e., the energy of latent representation concentrates in several dominant channels), and further observe two characteristics of nonlinear transforms to realize energy compaction (i.e., feature decorrelation and uneven energy modulation). However, they ignore to consider the evolution of energy compaction during training. In this paper, we further investigate this problem, and reveal that LIC models struggle to improve the efficiency of energy compaction through the training process. As demonstrated in Figure 1(a) (blue line), during training, an LIC model could use only 10% channels of the latent representation to preserve 87% of its energy in the first 0.5 million iterations, but would require extra 1.5 million iterations to further increase the energy by 6%. This suggests the necessity to ease energy compaction in training nonlinear transform and thus improve training efficiency.
@@ 3 Energy Based Interpretation of LIC Training
Understanding the training process of LIC is crucial to addressing the slow convergence issue. In this section, we begin by re-examining the results of previous work on the energy-compaction property of nonlinear transforms (Duan et al., 2022; He et al., 2022; Li et al., 2023), and then show the two reasons why the energy-compaction property is hard to learn.
Traditional transform coding usually leverages orthogonal linear transforms such as Discrete Cosine Transform (DCT), Discrete Wavelet Transform (DWT), Karhunen-Loève Transform (KLT), and multiscale geometric analysis to aggregate most energy of the image into few components. And the energy-compaction property has a significant contribution to the compression performance. In an optimized LIC model, the majority of energy in the latent representation is concentrated in a few dominant channels, while the other channels contain considerably less energy (as illustrated in Figure 2). According to (Li et al., 2023), the dominant channels holding high energy are low-frequency channels, whereas the others are high-frequency channels.
@@
Feature Decorrelation. Traditional transform coding inherently obtains energy-compacting representation through orthogonal linear transforms that are well designed for decorrelating frequency components in a structured manner. LIC learns nonlinear transforms to adapt to the source distribution of natural images and also achieves decorrelation in latent representation. In Figure 3, we estimate the average similarity between each two channels based on their guided backpropagation patterns (Springenberg et al., 2014) on the Kodak dataset throughout the whole training process. The pairwise channel similarity suggests the correlation of any two channels of latent representations is gradually reduced during the training process. However, unlike traditional transform coding, LIC requires numerous iterations in training to achieve ideal decorrelation, since orthogonality is not guaranteed for the analysis transform.
@@
Uneven Energy Modulation. Traditional transform coding commonly applies different quantization steps on the transform coefficients of different frequencies. The low-frequency coefficients usually adopt smaller quantization steps than high-frequency coefficients to reduce average distortion. On the contrary, LIC models employ uniform quantizers with a fixed quantization step of 1 for all the latent representation. Low-frequency channels that inherently carry higher energy require additional amplitude amplification (i.e., uneven energy modulation) to achieve finer quantization. Thus, the analysis transform gradually learns to amplify low-frequency channels, and overall energy in the latent representation increases as training proceeds. Figure 4 shows that the average intensity of latent representations is significantly scaled up from a value of 1 in the source space to over 100 in the latent space. However, the scaling-up primarily occurs in the last layer of neural network realizing the analysis transform. This causes unbalanced gradients that could slow down the training process.
@@ 4.2 Overall Architecture With Multi-stage Shortcuts
As illustrated in Figure 5, we apply four stacked WLS in the analysis transform and four stacked iWLS in the synthesis transform. For the analysis transform, the output of each WLS is concatenated with the output of each down-sampling layer of the nonlinear transform $g_{\textit{a}}$. The output of the last WLS (i.e., $\bm{\hat{P}}_{final}$) will be added to the output of analysis transform $\bm{F}$ to form the final latent representation $\bm{y}$ that is subsequently quantized and encoded. For the synthesis transform, iWLS implements the inverse operation of WLS in the analysis transform, where the DWT is replaced by the inverse DWT (IDWT) and the channel-wise multiplication in Equation 4 is replaced by a channel-wise division to achieve corresponding energy reduction.
@@ 5.1 Experimental Setup
Training. We apply our AuxT to several mainstream LIC models to show its effectiveness, including mb2018mean (Minnen et al., 2018), ELIC (He et al., 2022), STF (Zou et al., 2022), and TCM (Liu et al., 2023). All the models are trained on the ImageNet-1k (Deng et al., 2009) dataset and optimized using Adam optimizer (Kingma & Ba, 2014). We set the batch size to 16 for convolution-based LIC models (Minnen et al., 2018; He et al., 2022) and 8 for transformer-based LIC models Zou et al. (2022); Liu et al. (2023). We train the models without our AuxT for 0.6M and 2M iterations respectively, and train the models with our AuxT for 0.6M and 1M iterations, respectively. The learning rate is initialized as $10^{-4}$ and is decayed by a factor of 10 after 0.55M iterations for 0.6M training iterations scenario, after 0.9M iterations for 1M training iterations scenario, and after 1.8M iterations for 2M training iterations scenario.
@@ 5.4 Ablation Studies
Effect of the linearity of AuxT. Unlike the nonlinear transform used in LIC, our AuxT is a purely linear transform without any nonlinear operators. We demonstrate the necessity of this linearity by introducing a nonlinear layer after the OLP. As shown in Figure 8(d), adding either ReLU (Glorot et al., 2011) or GDN (Ballé et al., 2016a) significantly hinders the performance of AuxT. This is because the nonlinearity causes energy attenuation and disrupts orthogonality, which negatively impacts effective energy compaction. This demonstrates that a simple linear module can be more effective for energy compaction. More ablation studies can be found in Appendix A.3.
@@ A.1 Related Work
Learned Image Compression. Learned image compression has seen significant advancements in recent years due to its impressive rate-distortion (R-D) performance. Unlike traditional transform coding methods that use linear transforms to decorrelate input images, LIC employs nonlinear transforms to achieve a more compact latent representation. Early works (Ballé et al., 2016b; 2018; Minnen et al., 2018; Minnen & Singh, 2020) utilized convolutional neural networks (CNNs) to develop these nonlinear transforms, focusing on capturing local patterns in images. Later approaches incorporated non-local blocks (Cheng et al., 2020) to enhance the ability to capture global correlations. More recently, advancements have leveraged attention-based mechanisms and transformer architectures (Lu et al., 2022; Zou et al., 2022; Liu et al., 2023; Li et al., 2024a) for LIC, underscoring the idea that increasing complexity and capabilities can lead to more expressive nonlinear transforms.
Another key aspect of LIC research involves designing more effective entropy models. Earlier works (Ballé et al., 2016b; 2018) introduced factorized and hyperprior models to boost compression efficiency. More recent efforts (Minnen et al., 2018; Minnen & Singh, 2020; Koyuncu et al., 2022; Qian et al., 2022) have incorporated spatial or channel-wise autoregression into entropy models, further enhancing their performance.
@@
Energy Compaction Property in LIC. Energy compaction is a fundamental characteristic in traditional transform coding, indicating that most of the signal’s energy is concentrated into a few components. In the field of LIC, several studies have explored this concept. For instance, Cheng et al. (2019b) and Cheng et al. (2019a) introduced a spatial energy compaction-based penalty in the loss function, encouraging energy concentration within a few channels. He et al. (2022) observed that the analysis transform in LIC inherently exhibits an information compaction property, where a limited number of channels carry significantly more average energy. Additionally, Li et al. (2023) further demonstrated that higher-energy channels are typically characterized by low-frequency components, while lower-energy channels tend to carry high-frequency information. In our work, we find that the energy compaction process is closely linked to the slow training issue of LIC and propose an auxiliary transform to address this issue.
Decorrelation for LIC. Existing LIC methods often struggle with achieving better decorrelation ability to further remove the redundancies in the latent representation. The main solution is to leverage more complex neural network modules to enhance the overall representation ability of nonlinear transforms (Ballé et al., 2016b; 2018; Minnen et al., 2018; Minnen & Singh, 2020; Lu et al., 2022; Zou et al., 2022; Liu et al., 2023; Li et al., 2024a). In addition, Ballé et al. (2016a) proposed the Generalized Divisive Normalization (GDN) layer to helps to decorrelate information across channels. Guo et al. (2021) further explored the cross-channel relationships of the latents and achieve better channel-wise decorrelation by the proposed causal context prediction module.
@@
One key approach to speeding up training is the development of advanced optimizers. In addition to widely-used adaptive learning rate methods like Adam (Kingma & Ba, 2014) and accelerated schemes such as Nesterov momentum Nesterov (1983), several recent approaches have emerged to further enhance training efficiency. For example, Goyal (2017) introduced a scaling rule to adjust the learning rate, enabling faster training when using large mini-batches. Furthermore, You et al. (2017) and You et al. (2019) proposed layer-wise adaptive learning rates, which allow for the scaling of batch size, ultimately reducing training time.
@@
In end-to-end learned image compression, many works have introduced DWT into the nonlinear transform to enhance its decorrelation ability. Mishra et al. (2020) use DWT to preprocess the input image and separately encode its wavelet subbands. Ma et al. (2020) design a wavelet-like nonlinear transform using the lifting scheme. Fu et al. (2024) replace the down-sampling and up-sampling layers of LIC with DWT and IDWT, respectively, to better remove frequency-domain correlation. We highlight that our method differs from these methods as we do not modify the nonlinear transform itself, but instead introduce a bypass auxiliary transform to address the convergence issue. To further demonstrate the superiority of our method, we conduct experiments to test whether adopting DWT directly in the nonlinear transform can accelerate convergence. Specifically, we replace the sub-sampling and up-sampling layers of TCM with DWT and IDWT, respectively and append a Residual Block before and after the DWT and IDWT layers to maintain the parameter count.
@@ A.4 Energy Modulation Behavior
The orthogonality concept has been widely explored in the filed of deep learning (Le et al., 2011; Vorontsov et al., 2017; Bansal et al., 2018; Wang et al., 2020), as it implies energy preservation and encourages filter diversity. In our work, we use the orthogonal kernel regularization to force our orthogonal linear projection $\bm{W}\in\mathbb{R}^{4C\times D}$ as orthogonal layer. Specifically, if $4C\geq D$, $\bm{W}$ is an undercomplete matrix, and we use a row orthogonal regularizer $||\bm{W}^{T}\bm{W}-\bm{I}||_{F}^{2}$, where $\bm{I}\in\mathbb{R}^{D\times D}$ is the identify matrix. Conversely, if $4C\geq D$, $\bm{W}$ becomes an overcomplete matrix, and column orthogonality is encouraged using the penalty $||\bm{W}\bm{W}^{T}-\bm{I}^{\prime}||_{F}^{2}$. In our practical implementation, we use the row orthogonal regularizer for both cases, since the row orthogonality and column orthogonality are equivalent in the mean squared error (MSE).

### v3

@@ 1 Introduction
In fact, efficient training of LIC models has not been thoroughly explored, despite the properties of nonlinear transforms have been studied. Previous studies (Duan et al., 2022; He et al., 2022; Li et al., 2024c) find that the analysis transform in LIC exhibits energy compaction property (i.e., the energy of latent representation concentrates in several dominant channels), and further observe two characteristics of nonlinear transforms to realize energy compaction (i.e., feature decorrelation and uneven energy modulation). However, they ignore to consider the evolution of energy compaction during training. In this paper, we further investigate this problem, and reveal that LIC models struggle to improve the efficiency of energy compaction through the training process. As demonstrated in Figure 1(a) (blue line), during training, an LIC model could use only 10% channels of the latent representation to preserve 87% of its energy in the first 0.5 million iterations, but would require extra 1.5 million iterations to further increase the energy by 6%. This suggests the necessity to ease energy compaction in training nonlinear transform and thus improve training efficiency.
@@ 3 Energy Based Interpretation of LIC Training
Understanding the training process of LIC is crucial to addressing the slow convergence issue. In this section, we begin by re-examining the results of previous work on the energy-compaction property of nonlinear transforms (Duan et al., 2022; He et al., 2022; Li et al., 2024c), and then show the two reasons why the energy-compaction property is hard to learn.
Traditional transform coding usually leverages orthogonal linear transforms such as Discrete Cosine Transform (DCT), Discrete Wavelet Transform (DWT), Karhunen-Loève Transform (KLT), and multiscale geometric analysis to aggregate most energy of the image into few components. And the energy-compaction property has a significant contribution to the compression performance. In an optimized LIC model, the majority of energy in the latent representation is concentrated in a few dominant channels, while the other channels contain considerably less energy (as illustrated in Figure 2). According to (Li et al., 2024c), the dominant channels holding high energy are low-frequency channels, whereas the others are high-frequency channels.
@@
Feature Decorrelation. Traditional transform coding inherently obtains energy-compacting representation through orthogonal linear transforms that are well designed for decorrelating frequency components in a structured manner. LIC learns nonlinear transforms to adapt to the source distribution of natural images and also achieves decorrelation in latent representation. In Figure 3, we estimate the average similarity between each two channels based on their guided backpropagation patterns (Springenberg et al., 2015) on the Kodak dataset throughout the whole training process. The pairwise channel similarity suggests the correlation of any two channels of latent representations is gradually reduced during the training process. However, unlike traditional transform coding, LIC requires numerous iterations in training to achieve ideal decorrelation, since orthogonality is not guaranteed for the analysis transform.
@@
Uneven Energy Modulation. Traditional transform coding commonly applies different quantization steps on the transform coefficients of different frequencies. The low-frequency coefficients usually adopt smaller quantization steps than high-frequency coefficients to reduce average distortion. On the contrary, LIC models employ uniform quantizers with a fixed quantization step of 1 for all the latent representation. Low-frequency channels that inherently carry higher energy require additional amplitude amplification (i.e., uneven energy modulation) to achieve finer quantization. Thus, the analysis transform gradually learns to amplify low-frequency channels, and overall energy in the latent representation increases as training proceeds. Figure 4 shows that the max intensity of latent representations is significantly scaled up from a value of 1 in the source space to over 100 in the latent space. However, the scaling-up primarily occurs in the last layer of neural network realizing the analysis transform. This causes unbalanced gradients that could slow down the training process.
@@ 4.2 Overall Architecture With Multi-stage Shortcuts
As illustrated in Figure 5, we apply four stacked WLS in the analysis transform and four stacked iWLS in the synthesis transform. For the analysis transform, the output of each WLS will be added to the output of each down-sampling layer of the nonlinear transform $g_{\textit{a}}$. The output of the last WLS (i.e., $\bm{\hat{P}}_{final}$) will be added to the output of analysis transform $\bm{F}$ to form the final latent representation $\bm{y}$ that is subsequently quantized and encoded. For the synthesis transform, iWLS implements the inverse operation of WLS in the analysis transform, where the DWT is replaced by the inverse DWT (IDWT) and the channel-wise multiplication in Equation 4 is replaced by a channel-wise division to achieve corresponding energy reduction.
@@ 5.1 Experimental Setup
Training. We apply our AuxT to several mainstream LIC models to show its effectiveness, including mb2018mean (Minnen et al., 2018), ELIC (He et al., 2022), STF (Zou et al., 2022), and TCM (Liu et al., 2023). All the models are trained on the ImageNet-1k (Deng et al., 2009) dataset and optimized using Adam optimizer (Kingma & Ba, 2015). We set the batch size to 16 for convolution-based LIC models (Minnen et al., 2018; He et al., 2022) and 8 for transformer-based LIC models Zou et al. (2022); Liu et al. (2023). We train the models without our AuxT for 0.6M and 2M iterations respectively, and train the models with our AuxT for 0.6M and 1M iterations, respectively. The learning rate is initialized as $10^{-4}$ and is decayed by a factor of 10 after 0.55M iterations for 0.6M training iterations scenario, after 0.9M iterations for 1M training iterations scenario, and after 1.8M iterations for 2M training iterations scenario.
@@ 5.4 Ablation Studies
Effect of the linearity of AuxT. Unlike the nonlinear transform used in LIC, our AuxT is a purely linear transform without any nonlinear operators. We demonstrate the necessity of this linearity by introducing a nonlinear layer after the OLP. As shown in Figure 8(d), adding either ReLU (Glorot et al., 2011) or GDN (Ballé et al., 2016) significantly hinders the performance of AuxT. This is because the nonlinearity causes energy attenuation and disrupts orthogonality, which negatively impacts effective energy compaction. This demonstrates that a simple linear module can be more effective for energy compaction. More ablation studies can be found in Appendix A.3.
@@ 6 Conclusion
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grant 62320106003, Grant 62371288, Grant 62301299, Grant 62401357, Grant 62431017, Grant 62401366, Grant 62120106007, Grant 62125109, Grant U24A20251, and in part by the Program of Shanghai Science and Technology Innovation Project under Grant 24BC3200800.
@@ A.1 Related Work
Learned Image Compression. Learned image compression has seen significant advancements in recent years due to its impressive rate-distortion (R-D) performance. Unlike traditional transform coding methods that use linear transforms to decorrelate input images, LIC employs nonlinear transforms to achieve a more compact latent representation. Early works (Ballé et al., 2017; 2018; Minnen et al., 2018; Minnen & Singh, 2020) utilized convolutional neural networks (CNNs) to develop these nonlinear transforms, focusing on capturing local patterns in images. Later approaches incorporated non-local blocks (Cheng et al., 2020) to enhance the ability to capture global correlations. More recently, advancements have leveraged attention-based mechanisms and transformer architectures (Lu et al., 2022; Zou et al., 2022; Liu et al., 2023; Li et al., 2024a) for LIC, underscoring the idea that increasing complexity and capabilities can lead to more expressive nonlinear transforms.
Another key aspect of LIC research involves designing more effective entropy models. Earlier works (Ballé et al., 2017; 2018) introduced factorized and hyperprior models to boost compression efficiency. More recent efforts (Minnen et al., 2018; Minnen & Singh, 2020; Koyuncu et al., 2022; Qian et al., 2022) have incorporated spatial or channel-wise autoregression into entropy models, further enhancing their performance.
@@
Energy Compaction Property in LIC. Energy compaction is a fundamental characteristic in traditional transform coding, indicating that most of the signal’s energy is concentrated into a few components. In the field of LIC, several studies have explored this concept. For instance, Cheng et al. (2019b) and Cheng et al. (2019a) introduced a spatial energy compaction-based penalty in the loss function, encouraging energy concentration within a few channels. He et al. (2022) observed that the analysis transform in LIC inherently exhibits an information compaction property, where a limited number of channels carry significantly more average energy. Additionally, Li et al. (2024c) further demonstrated that higher-energy channels are typically characterized by low-frequency components, while lower-energy channels tend to carry high-frequency information. In our work, we find that the energy compaction process is closely linked to the slow training issue of LIC and propose an auxiliary transform to address this issue.
Decorrelation for LIC. Existing LIC methods often struggle with achieving better decorrelation ability to further remove the redundancies in the latent representation. The main solution is to leverage more complex neural network modules to enhance the overall representation ability of nonlinear transforms (Ballé et al., 2017; 2018; Minnen et al., 2018; Minnen & Singh, 2020; Lu et al., 2022; Zou et al., 2022; Liu et al., 2023; Li et al., 2024a). In addition, Ballé et al. (2016) proposed the Generalized Divisive Normalization (GDN) layer to helps to decorrelate information across channels. Guo et al. (2021) further explored the cross-channel relationships of the latents and achieve better channel-wise decorrelation by the proposed causal context prediction module.
@@
One key approach to speeding up training is the development of advanced optimizers. In addition to widely-used adaptive learning rate methods like Adam (Kingma & Ba, 2015) and accelerated schemes such as Nesterov momentum Nesterov (1983), several recent approaches have emerged to further enhance training efficiency. For example, Goyal et al. (2017) introduced a scaling rule to adjust the learning rate, enabling faster training when using large mini-batches. Furthermore, You et al. (2017) and You et al. (2020) proposed layer-wise adaptive learning rates, which allow for the scaling of batch size, ultimately reducing training time.
@@
In end-to-end learned image compression, many works have introduced DWT into the nonlinear transform to enhance its decorrelation ability. Mishra et al. (2020) use DWT to preprocess the input image and separately encode its wavelet subbands. Ma et al. (2020) design a wavelet-like nonlinear transform using the lifting scheme. Fu et al. (2024) replace the final down-sampling and up-sampling layers of LIC with DWT and IDWT, respectively, to better remove frequency-domain correlation. We highlight that our method differs from these methods as we do not modify the nonlinear transform itself, but instead introduce a bypass auxiliary transform to address the convergence issue. To further demonstrate the superiority of our method, we conduct experiments to evaluate whether adopting DWT directly in the nonlinear transform can accelerate convergence. Specifically, we replace the sub-sampling and up-sampling layers of TCM with DWT and IDWT, respectively and append a Residual Block before and after the DWT and IDWT layers to maintain the parameter count.
@@ A.4 Energy Modulation Behavior
The orthogonality concept has been widely explored in the filed of deep learning (Le et al., 2011; Vorontsov et al., 2017; Bansal et al., 2018; Wang et al., 2020), as it implies energy preservation and encourages filter diversity. In our work, we use the orthogonal kernel regularization to force our orthogonal linear projection $\bm{W}\in\mathbb{R}^{4C\times D}$ as orthogonal layer. Specifically, if $4C>D$, $\bm{W}$ is an undercomplete matrix, and we use a row orthogonal regularizer $||\bm{W}^{T}\bm{W}-\bm{I}||_{F}^{2}$, where $\bm{I}\in\mathbb{R}^{D\times D}$ is the identify matrix. Conversely, if $4C<D$, $\bm{W}$ becomes an overcomplete matrix, and column orthogonality is encouraged using the penalty $||\bm{W}\bm{W}^{T}-\bm{I}^{\prime}||_{F}^{2}$. In our practical implementation, we use the row orthogonal regularizer for both cases, since the row orthogonality and column orthogonality are equivalent in the mean squared error (MSE).